OpenStudy (anonymous):

Explain how to use prime factorization to find the LCM of two numbers. Include an example.

OpenStudy (anonymous):

like \[4x ^{2}-3x\]

OpenStudy (anonymous):

i actually tried that a few times and it didnt work srrrry

OpenStudy (anonymous):

suppose you have two numbers like 12 and 80 12=2*6 12=2*2*3 80=2*40 80=2*2*20 80=2*2*2*10 80=2*2*2*2*5 80 has the highest number of 2's, 4 f them, it also has a five 12 has a three not covered by the 80 so a multiple of both 12 and 80 has to be divisible by 2 four times by five and by three . we can multiply those together to get the lowest multiple 2*2*2*2*3*5=240 so 240 is the lowest common multiple